8. A comittee of 10 women and 9 men is looking to create a 6 person committee.
a. how many ways are there of choosing this committee?
19*18*17*16*15*14=19535040
b. How many ways are there of chosing 4 men and 2 women?
(i'm confused)
c. Whats the probability that it's selected with 4 men and 2 women?
(I havent a clue)
d. What is the probility that it's selected with 5 women and 1 man?
( i do not know)
Please help! Thanks.

Found 2 solutions by richard1234, stanbon:
Answer by richard1234(7193)

a: Order doesn't matter, so it is 19C6 = (19*18*...*14)/6!

b: Choose the men separately, and the women separately. Hence there are 9C4*10C2 ways.

c: The probability is simply the answer to b) over the answer to a), assuming all combinations are equally likely.

d: We do d) the same way as we did b). There are 10C5*9C1 ways to select such a committee. Divide this by the total number to obtain the probability.

Answer by stanbon(75887)

. A comittee of 10 women and 9 men is looking to create a 6 person committee.
a. how many ways are there of choosing this committee?
---
Ans: 19C6 = 19!/(13!*6!) = (19*18*...*14)/(1*2*...6) = 27,132
---------------------------------------
b. How many ways are there of chosing 4 men and 2 women?
---
Ans: [9C4*10C2] = 5670
---------------------------------------
c. Whats the probability that it's selected with 4 men and 2 women?
Ans: 5670/27132 = 0.2090
=================================
d. What is the probility that it's selected with 5 women and 1 man?
Ans: [10C5*9C1]/[19C5] = 0.0836
===================================
Cheers,
Stan H.